# Option Pricing with Time-Varying Volatility Risk Aversion

### v3

@@ 5.5 Key Insight from Autocorrelations of Pricing Error
Figure 7: The autocorrelation function (ACF) for derivative pricing errors over the full sample period (1990-2022). The ACFs for the VIX are shown in the upper panels, and the ACFs for the average option pricing errors are shown in the lower panels. The left panels present the results for the CHNG model, while the right panels present the results for the DHNG model.
@@ 5.6 Adaptive Compensation for Misspecification
As a robustness check, we computed an alternative VRR that does not rely on the Heston-Nandi GARCH model. This alternative measure is based on model-free realized variances and uses a simple AR(1) model to define expected variance under $\mathbb{P}$. The resulting time series of this alternative VRR, reported in Appendix B, closely resembles that of $\eta_{t}$ in Figure 5. This similarity suggests that the observed fluctuations in $\eta_{t}$ are not driven by a flaw specific to the GARCH model (1)-(2). However, an oversimplified description of how expected volatility evolves can be the cause. The AR(1) model and standard GARCH models both produce volatility forecasts with mean-reverting characteristics.
Regarding misspecification as a potential source of time variation in $\eta_{t}$, we generated scatterplots of pricing errors against $\eta_{t}$. These scatterplots do not suggest a systematic relationship between pricing errors and $\eta_{t}$; see Figure B.2 in Appendix B.
@@ 6 Variance Risk Aversion and Economic Fundamentals
We also include the variance risk premium (VRP) by Carr and Wu, (2008), which Grith et al., (2017) suggest serves as a proxy for market uncertainty. Additionally, we incorporate three uncertainty measures from existing literature: the Economic Uncertainty Index by Bali et al., (2014), the Survey-based Uncertainty Index by Ozturk and Sheng, (2018); [^20] This index is constructed from the Consensus Forecasts publication by Consensus Economics Inc. and the Sentiment variable by Baker and Wurgler, (2006).
@@
Economic Policy Uncertainty 0.258*** 0.203***
@@
Survey-based Uncertainty 0.106*** 0.065***
@@
Economic Uncertainty Index 0.037*** -0.007
@@
Variance Risk Premium 0.097*** 0.074***
@@
Stock Market Volatility 0.317*** 0.318*** 0.299*** 0.300*** 0.239*** 0.297*** 0.290*** 0.202*** 0.148***
@@
Default Yield Spread 0.162*** 0.108** 0.132*** 0.143*** 0.198*** 0.093* 0.080 0.160*** 0.100**
@@
Default Return Spread 0.031*** 0.031*** 0.029*** 0.028*** 0.026*** 0.029*** 0.027*** 0.024*** 0.021***
@@
Long Term Yield 0.040*** 0.046*** 0.027*** 0.027*** 0.056*** 0.021** 0.035*** 0.033*** 0.033***
@@
Dividend Price Ratio -0.158** -0.158** -0.118* -0.135* -0.233*** -0.116* -0.165** -0.174*** -0.179***
@@
$R^{2}$ 0.649 0.693 0.681 0.663 0.683 0.667 0.670 0.770 0.829
$\rho(1)$ 0.617 0.566 0.581 0.611 0.607 0.625 0.628 0.387 0.368
Note: The control variables are from Welch and Goyal, (2007) and Baker and Wurgler, (2006). “Other Controls” refer to six macro variables that control for macroeconomic conditions (monthly growth rate of industrial production, real durables consumption, real nondurables consumption, real services consumption, employment, and the NBER recession indicator) and five financial variables (earnings-to-price ratio, book-to-market ratio, net equity expansion, long-term rate of return, and the consumer price index). Most of these additional control variables are insignificant.
The regression results are presented in Table 6. The seven core variables are listed at the top of the table, followed by the six control variables that contributed the most to explaining variation in $\eta_{t}$. There are eleven additional control variables, labeled “Other Controls”, which were largely insignificant. [^21] We follow Welch and Goyal, (2007) in the definitions of these variables. For instance, the yield spread is the difference between BAA- and AAA-rated corporate bond yields, and the default return spread is the difference between the return on long-term corporate bonds and government bonds. See Welch and Goyal, (2007) for more details. To conserve space, the results for all control variables (along with their labels) are presented in the Appendix B.
The first column in Table 6 corresponds to the specification that includes only the control variables, which explain 64.9% of the variation in the monthly average of $\eta$. The six most significant control variables are stock market volatility, the dividend-price ratio, and four variables related to interest rates and credit risk.
@@
Finally, the variance risk premium is an empirical measure of the difference between volatility under $\mathbb{Q}$ and $\mathbb{P}$. As expected, this variable has a positive coefficient and contributes the most to the increase in $R^{2}$. The last column shows the results for a “kitchen sink regression” that includes all explanatory variables. Most of the variables in Table 6 remain significant in this regression, and the $R^{2}$ increases to nearly 83%. However, two of the core variables, Forecast Dispersion of GDP Growth and the Economic Uncertainty Index, become insignificant in the kitchen sink regression, possibly due to collinearity, as these two variables have the highest correlation (65.7%) among all the core variables.
@@ A.1 Proof of Theorem 1
$=$ $\exp\left[v_{1}r+v_{2}\omega+[v_{1}(\lambda-\tfrac{1}{2})+v_{2}(\beta+\alpha\gamma^{2})]h_{t+1}\right]\times\text{$\mathbb{E}_{t}^{\mathbb{P}}$}\left[v_{2}\alpha z_{t+1}^{2}+(v_{1}-2v_{2}\alpha\gamma)\sqrt{h_{t+1}}z_{t+1}\right]$
@@
$z_{t+1}^{*}=\frac{z_{t+1}-\mathbb{E}_{t}^{\mathbb{Q}}(z_{t+1})}{\sqrt{\mathrm{var}_{t}^{\mathbb{Q}}(z_{t+1})}}=\frac{1}{\sqrt{\eta_{t}}}\left(z_{t+1}+\left(\lambda+\frac{1}{2}\eta_{t}-\tfrac{1}{2}\right)\sqrt{h_{t+1}}\right),\quad t=1,2,\ldots,$ (A.5)
such that $z_{t}^{\ast}\sim iid\ N(0,1)$ under $\mathbb{\mathbb{Q}}$. Substituting the risk-neutralized quantities, $z_{t+1}^{*}$ and $h_{t+1}^{*}=\eta_{t}h_{t+1}$ into (1), we arrive at the expressions stated in part $(ii)$. Finally, part $(iii)$ is an implication of the MGF for $z_{t+1}$ under $\mathbb{\mathbb{Q}}$, see (A.4). $\square$
A.2 Proof of Equation (10)
@@
$=\tfrac{1}{2}\left(1-2\alpha\xi_{t}\right)\left(\lambda-\tfrac{1}{2}\right)^{2}-\tfrac{1}{2}\left(\lambda-\tfrac{1}{2}\right)^{2}+\lambda\left(\lambda-\tfrac{1}{2}\right)-\tfrac{1}{2}\left[\tfrac{1}{\eta_{t}}\left(\lambda-\tfrac{1}{2}\right)^{2}+\eta_{t}+(\lambda-\tfrac{1}{2})\right]$
$=-\tfrac{1}{2}\eta_{t}+\tfrac{1}{2}\left(\lambda-\tfrac{1}{2}\right)^{2}.$
@@ A.3 Proof of Theorem 2
where $\nu_{t}=\alpha_{t}^{*}((z_{t}^{*})^{2}-1-2\gamma_{t-1}^{*}z_{t}^{*}\sqrt{h_{t}^{*}})$ is a zero-mean variation condition on $\mathcal{F}_{t-1}$. We denote $A_{t}\equiv\omega_{t}+\alpha_{t}$ and $B_{t}\equiv\beta_{t}^{\ast}+\alpha_{t}\gamma_{t-1}^{\ast 2}$, then we have
@@
$+\sum_{i=1}^{t+k-1}B_{t+k-2}\cdots B_{i}\nu_{i}.$
In order to get its analytical formula, we replace $B_{t}$ with $\tilde{B}_{t}\equiv\beta_{t}^{\ast}+\alpha_{t}\tilde{\gamma}_{t-1}^{\ast 2}$ where $\tilde{\gamma}_{t}^{\ast}=\tfrac{1}{\eta_{t}}(\gamma+\lambda)$. Empirically $\tilde{\gamma}_{t}^{\ast}$ is indistinguishable from $\gamma_{t}^{*}=\tfrac{1}{\eta_{t}}(\gamma+\lambda+\tfrac{1}{2}(\eta_{t}-1))$ because the magnitude of $\gamma+\lambda$ is much larger than $\frac{1}{2}(\eta_{t}-1)$. Empirically we have $\gamma+\lambda\approx 280$ whereas $\frac{1}{2}(\eta_{t}-1)$ ranges between $-0.25$ and 1.5, with an average value of about 0.1.
It follows that [^22] Recall that $\alpha_{t}^{\ast}=\alpha\eta_{t}\eta_{t-1}$ and $\beta_{t}^{*}=\beta\tfrac{\eta_{t}}{\eta_{t-1}}$.
@@
Suppose that $\varepsilon_{t}$ is iid and let $\Psi(\phi)\equiv\log\mathbb{E}[\exp(\phi\varepsilon_{t})]$ denote its log-MGF. Then we have
@@
$\mathbb{E}_{t}\left(\eta_{t+k}\right)=\exp\left\{(1-\varphi^{k})\zeta+\varphi^{k}\log\eta_{t}+\sum_{j=1}^{k}\Psi(\varphi^{j})\right\}$, and
@@
$\Lambda_{1}(k)$ $=\exp\left\{\left(1-\varphi^{k-1}\right)\zeta+\sum_{j=1}^{k-1}\Psi(\varphi^{j})\right\},$
@@
If $\varepsilon_{t}$ is normally distributed with variance $\sigma^{2}$, we have $\Psi(\phi)=\frac{1}{2}\sigma^{2}\phi^{2}$. $\square$
@@
We have $B_{t}=\beta_{t}^{\ast}+\alpha_{t}^{\ast}\gamma_{t-1}^{\ast 2}\in[\tilde{\beta}_{{\rm min}},\tilde{\beta}_{{\rm max}}]\frac{\eta_{t}}{\eta_{t-1}}$. Since $\delta(x)$ is a monotone function in $x$, there exists an $\bar{\beta}\in[\tilde{\beta}_{{\rm min}},\tilde{\beta}_{{\rm max}}]$, such that $\delta(\bar{\beta})=A\times\sqrt{\tfrac{1}{M}\sum_{k=1}^{M}\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+k}^{*})}$ (i.e., true value of VIX), and $\Delta^{{\rm vix}}=|\delta(\bar{\beta})-\delta(\tilde{\beta})|$, then we have $\Delta\leq\delta(\tilde{\beta}_{{\rm max}})-\delta(\tilde{\beta}_{{\rm min}})$. Note that $\delta(x)$ can be expressed in terms of $a_{1}(\eta_{t},M,\sigma^{2})$ and $a_{2}(\eta_{t},M,\sigma^{2})$. ∎
@@ A.4 Proof of Lemma 2
The appropriate MGF for cumulative returns, for the case where $\eta_{t}$ is predetermined is denoted $g_{t,M}(s|\bar{\boldsymbol{\eta}}_{t,M})$ in the following Lemma A.2.
Lemma A.2.
In the predetermined case, the MGF for cumulative returns has the exponentially affine form:
@@
For later use, we note that the MGF for $(R_{t+1},h_{t+2}^{*})$ is given by:

### v4

@@ 4.5 Key Insight from Autocorrelations of Pricing Error
Figure 7: The autocorrelation function (ACF) for derivative pricing errors over the full sample period (January 1990 to December 2021). The ACFs for the VIX are shown in the upper panels, and the ACFs for the average option pricing errors are shown in the lower panels. The CHNG and DHNG models are estimated with VIX data in the upper panels and option prices in the lower panels. Model parameters are based on the Table 3.
Alt text: Autocorrelation functions (ACFs) of derivative pricing errors over the full sample period (1990–2021) for the CHNG and DHNG models. DHNG exhibits vastly smaller autocorrelations for both VIX pricing errors and average option pricing errors, consistent with improved pricing accuracy.
@@ 4.6 Adaptive Compensation for Misspecification
As a robustness check, we computed an alternative VRR that does not rely on the Heston-Nandi GARCH model. This alternative measure is based on model-free realized variances and uses a simple AR(1) model to define expected variance under $\mathbb{P}$. The resulting time series of this alternative VRR, reported in Appendix S.2, closely resembles that of $\eta_{t}$ in Figure 5. This similarity suggests that the observed fluctuations in $\eta_{t}$ are not driven by a flaw specific to the GARCH model (1)-(2). However, an oversimplified description of how expected volatility evolves can be the cause. The AR(1) model and standard GARCH models both produce volatility forecasts with mean-reverting characteristics.
Regarding misspecification as a potential source of time variation in $\eta_{t}$, we generated scatterplots of pricing errors against $\eta_{t}$. These scatterplots do not suggest a systematic relationship between pricing errors and $\eta_{t}$; see Figure S.2 in Appendix S.2.
@@ 5 Variance Risk Aversion and Economic Fundamentals
We also include the variance risk premium (VRP) by Carr and Wu, (2008), which Grith et al., (2017) suggest serves as a proxy for market uncertainty. Additionally, we incorporate three uncertainty measures from existing literature: the Economic Uncertainty Index by Bali et al., (2014), the Survey-based Uncertainty Index by Ozturk and Sheng, (2018); [^21] This index is constructed from the Consensus Forecasts publication by Consensus Economics Inc. The authors provide data up to October 2021, and we fill in the values for the last two months using a simple predictive regression on all our explanatory variables. and the well-known Sentiment variable by Baker and Wurgler, (2006).
@@
Economic Policy Uncertainty 0.380*** 0.208***
@@
Survey-based Uncertainty 0.148*** 0.078***
@@
Economic Uncertainty Index 0.071*** 0.013
@@
Variance Risk Premium 0.120*** 0.096***
@@
Stock Market Volatility 0.394*** 0.395*** 0.368*** 0.356*** 0.294*** 0.363*** 0.354*** 0.395*** 0.306***
@@
Default Yield Spread 0.385*** 0.358*** 0.372*** 0.365*** 0.393*** 0.282*** 0.246*** 0.285*** 0.196***
@@
Default Return Spread 0.026*** 0.027*** 0.025*** 0.025*** 0.019*** 0.024*** 0.021*** 0.026*** 0.020***
@@
Long Term Yield 0.028** 0.030** 0.040*** 0.043*** 0.077*** 0.016 0.045*** 0.035*** 0.063***
@@
Dividend Price Ratio -0.457*** -0.435*** -0.525*** -0.529*** -0.638*** -0.463*** -0.403*** -0.354*** -0.471***
@@
$R^{2}$ 0.709 0.722 0.722 0.732 0.756 0.741 0.761 0.817 0.866
$\rho(1)$ 0.607 0.579 0.620 0.628 0.612 0.626 0.584 0.326 0.383
Note: The full-sample monthly regression results from January 1990 to December 2021, and the dependent variable is logarithm of monthly averaged variance risk ratio estimated by option prices (the last column of Table 3). The control variables are from Welch and Goyal, (2007) and Baker and Wurgler, (2006). “Other Controls” refer to six macro variables that control for macroeconomic conditions (monthly growth rate of industrial production, real durables consumption, real nondurables consumption, real services consumption, employment, and the NBER recession indicator) and five financial variables (earnings-to-price ratio, book-to-market ratio, net equity expansion, long-term rate of return, and the consumer price index). Most of these additional control variables are insignificant. Newey-West standard errors are reported in parentheses. $\rho(1)$ the first-order autocorrelation for residuals.
The full-sample monthly regression results are presented in Table 6. The seven core variables are listed at the top of the table, followed by the six control variables that contributed the most to explaining variation in $\eta_{t}$. There are eleven additional control variables, labeled “Other Controls”, which were largely insignificant. [^22] We follow Welch and Goyal, (2007) in the definitions of these variables. For instance, the yield spread is the difference between BAA- and AAA-rated corporate bond yields, and the default return spread is the difference between the return on long-term corporate bonds and government bonds. See Welch and Goyal, (2007) for more details. To conserve space, the results for all control variables (along with their labels) are presented in the Appendix S.4.
The first column in Table 6 corresponds to the specification that includes only the control variables, which explain 70.9% of the variation in the monthly average of $\eta$. The six most significant control variables are stock market volatility, the dividend-price ratio, and four variables related to interest rates and credit risk.
@@
Finally, the variance risk premium is an empirical measure of the difference between volatility under $\mathbb{Q}$ and $\mathbb{P}$. As expected, this variable has a positive coefficient and contributes the most to the increase in $R^{2}$. The last column shows the results for a “kitchen sink regression” that includes all explanatory variables. Most of the variables in Table 6 remain significant in this regression, and the $R^{2}$ increases to nearly 87%. However, three core variables–Forecast Dispersion of Unemployment, Forecast Dispersion of GDP Growth, and the Economic Uncertainty Index–become insignificant in the kitchen sink regression, possibly due to collinearity, as the first two and the last two variables exhibit the highest correlations among all core variables (88.6% and 54.8%), respectively.
@@ A.1 Proof of Theorem 1
$=$ $\exp\left[v_{1}r+v_{2}\omega+[v_{1}(\lambda-\tfrac{1}{2})+v_{2}(\beta+\alpha\gamma^{2})]h_{t+1}\right]\times\text{$\mathbb{E}_{t}^{\mathbb{P}}$}\left[\exp\left(v_{2}\alpha z_{t+1}^{2}+(v_{1}-2v_{2}\alpha\gamma)\sqrt{h_{t+1}}z_{t+1}\right)\right]$
@@
$z_{t+1}^{*}=\frac{z_{t+1}-\mathbb{E}_{t}^{\mathbb{Q}}(z_{t+1})}{\sqrt{\mathrm{var}_{t}^{\mathbb{Q}}(z_{t+1})}}=\frac{1}{\sqrt{\eta_{t}}}\left(z_{t+1}+\left(\lambda+\frac{1}{2}\eta_{t}-\tfrac{1}{2}\right)\sqrt{h_{t+1}}\right),$
such that $z_{t}^{\ast}|\mathcal{G}_{t}\sim iid\ N(0,1)$ under $\mathbb{\mathbb{Q}}$. Substituting the risk-neutralized quantities, $z_{t+1}^{*}$ and $h_{t+1}^{*}=\eta_{t}h_{t+1}$ into (1), we arrive at the expressions stated in part $(ii)$. Finally, part $(iii)$ is an implication of the MGF for $z_{t+1}$ under $\mathbb{\mathbb{Q}}$, see (A.4). $\square$
A.2 Proof of Lemma 1
@@
$=\tfrac{1}{2}\left(1-2\alpha\xi_{t}\right)\left(\lambda-\tfrac{1}{2}\right)^{2}-\tfrac{1}{2}\left(\lambda-\tfrac{1}{2}\right)^{2}+\lambda\left(\lambda-\tfrac{1}{2}\right)-\tfrac{1}{2}\left[\tfrac{1}{\eta_{t}}\left(\lambda-\tfrac{1}{2}\right)^{2}+\tfrac{1}{4}\eta_{t}+(\lambda-\tfrac{1}{2})\right]$
$=-\tfrac{1}{8}\eta_{t}+\tfrac{1}{2}\left(\lambda-\tfrac{1}{2}\right)^{2}.$
@@ A.3 Proof of Theorem 2
where $\nu_{t}=\alpha_{t}^{*}((z_{t}^{*})^{2}-1-2\gamma_{t-1}^{*}z_{t}^{*}\sqrt{h_{t}^{*}})$ is a zero-mean variation condition on $\mathcal{G}_{t-1}$ due to the facts that $z_{t}^{*}|\mathcal{G}_{t-1}\overset{\mathbb{Q}}{\sim}iid\ N\left(0,1\right)$, and the independence between $\varepsilon_{t}$ and $z_{t}$.
We denote $A_{t}\equiv\omega_{t}^{*}+\alpha_{t}^{*}$ and $B_{t}\equiv\beta_{t}^{\ast}+\alpha_{t}^{*}\gamma_{t-1}^{\ast 2}$, then we have
@@
$+\sum_{i=1}^{k-2}B_{t+k-1}B_{t+k-2}\cdots B_{t+i+1}\nu_{t+i}+\nu_{t+k-1}.$
In order to get its analytical formula, we replace $B_{t}$ with $\tilde{B}_{t}\equiv\beta_{t}^{\ast}+\alpha_{t}^{*}\tilde{\gamma}_{t-1}^{\ast 2}$ where $\tilde{\gamma}_{t}^{\ast}=\tfrac{1}{\eta_{t}}(\gamma+\lambda)$. Empirically $\tilde{\gamma}_{t}^{\ast}$ is indistinguishable from $\gamma_{t}^{*}=\tfrac{1}{\eta_{t}}(\gamma+\lambda+\tfrac{1}{2}(\eta_{t}-1))$ because the magnitude of $\gamma+\lambda$ is much larger than $\frac{1}{2}(\eta_{t}-1)$. Empirically we have $\gamma+\lambda\approx 253$ whereas $\frac{1}{2}(\eta_{t}-1)$ ranges between $-0.27$ and 1.39, with an average value of about 0.07. Then it follows that [^23] Recall that $\alpha_{t}^{\ast}=\alpha\eta_{t}\eta_{t-1}$ and $\beta_{t}^{*}=\beta\tfrac{\eta_{t}}{\eta_{t-1}}$.
@@
Suppose that $\varepsilon_{t}$ is iid and let $\Psi(s)\equiv\log\mathbb{E}[\exp(s\varepsilon_{t})]$ denote its log-MGF. Then we have
@@
$\mathbb{E}_{t}\left(\eta_{t+k}\right)=\exp\left\{(1-\varphi^{k})\zeta+\varphi^{k}\log\eta_{t}+\sum_{j=0}^{k-1}\Psi(\varphi^{j})\right\}$, and
@@
$\Lambda_{1}(k)$ $=\exp\left\{\left(1-\varphi^{k-1}\right)\zeta+\sum_{j=0}^{k-2}\Psi(\varphi^{j})\right\},$
@@
If $\varepsilon_{t}$ is normally distributed with zero mean and variance $\sigma^{2}$, we have $\Psi(\phi)=\frac{1}{2}\sigma^{2}\phi^{2}$. $\square$
@@
We have $B_{t}=\beta_{t}^{\ast}+\alpha_{t}^{\ast}\gamma_{t-1}^{\ast 2}\in[\tilde{\beta}_{{\rm min}},\tilde{\beta}_{{\rm max}}]\frac{\eta_{t}}{\eta_{t-1}}$. Since $\delta(x)$ is a monotone function in $x$, there exists an $\bar{\beta}\in[\tilde{\beta}_{{\rm min}},\tilde{\beta}_{{\rm max}}]$, such that $\delta(\bar{\beta})=A\times\sqrt{\tfrac{1}{M}\sum_{k=1}^{M}\mathbb{E}_{t}^{\mathbb{Q}}(h_{t+k}^{*})}$ (i.e., true value of VIX), and $\Delta=|\delta(\bar{\beta})-\delta(\tilde{\beta})|$, then we have $\Delta\leq\delta(\tilde{\beta}_{{\rm max}})-\delta(\tilde{\beta}_{{\rm min}})$. Note that $\delta(x)$ can be expressed in terms of $a_{1}(\eta_{t},M,\sigma^{2})$ and $a_{2}(\eta_{t},M,\sigma^{2})$. ∎
@@ A.4 Proof of Lemma 2
Lemma A.2 (Full Version of Lemma 2).
In the predetermined case with path $\bar{\boldsymbol{\eta}}_{t,M}$, the conditional MGF for future cumulative returns has the following exponentially affine form:
@@
For later use, we note that the MGF for $(R_{t+1},h_{t+2}^{*})$ is given by